write an equation in slope-intercept form of the line satisfying the given conditions
the line passes through (8,-3) and is parallel to the ling whose equation is y=2x+4

Answer by macston(5194)

y=mx+b To be parallel, slope=m must equal 2.
y=2x+b Plug in known point to solve for b.
-3=2(8)+b
-3=16+b
-19=b b=-19 so the final equation:
y=2x-19 ANSWER: The equation in slope-intercept form is y=2x-19.